In an experiment to determine the formula of an oxide of lead metal,excess hydrogen gas was passed over 100g of the heated oxide 8.67g lead metal was formed.calculate the formula of the oxide
1
Expert's answer
2020-04-08T13:26:56-0400

Remark:

Apparently the question contains a typo, and the correct mass of oxide is 10.0 g


So, the question:

In an experiment to determine the formula of an oxide of lead metal,excess hydrogen gas was passed over 10.0g of the heated oxide 8.67g lead metal was formed.calculate the formula of the oxide

Solution:

PbxOy+yH2 "\to" xPb+yH2O - general formula of reaction

mo=mPbxOy-mPb=10.0g-8.67g=1.33g

no=mo/Ao=1.33g/(16g*mol-1)=0.083mol

nPb=mPb/APb=8.67g/(207g*mol-1)=0.042mol

x/y=nPb/no=0.042mol/0.083mol =1/2

So, the formula of oxide is PbO2

Answer: PbO2